OpenStudy (anonymous):

about calculus ... 2[3-2(4-8)]=.....

OpenStudy (anonymous):

100 bucks says i'mma Medal Saifoo when he posts his answer!

OpenStudy (saifoo.khan):

Ninjád!

OpenStudy (saifoo.khan):

hehel.

OpenStudy (anonymous):

Damn! :( I tried to hurry too!!! haha

OpenStudy (saifoo.khan):

haha.

OpenStudy (anonymous):

I think easy.. for u... haha

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

22

OpenStudy (anonymous):

4-8=-4 -2*-4=8 3+8=11 2*11=22